X^2-2x-3=-3
THANK YOU

Answer by stanbon(75887)

X^2-2x-3=-3
------
x^2 - 2x = 0
----
Factor::
x (x-2) = 0
---
Ans: x = 0 or x = 2
===========
Cheers,
Stan H.
-------------